OpenStudy (insa):

Simplify the following giving your answer in positive indices a/b + (b/-a)^-1

OpenStudy (gunsodo):

\[\frac{ a }{ b }+(\frac{ b }{ -a})^-1\] =\[\frac{ a }{ b }+\frac{ -a }{ b }\] = 0